OpenStudy (tparker22):

If 14.3 moles of H2O2 is decomposed, how many grams of oxygen gas are produced? Show ALL your work to get full credit.

OpenStudy (whpalmer4):

This is really a question for the chemistry section, but as you've already posted it here... any thoughts on how to approach this?

OpenStudy (tparker22):

I just realized I posted it in the Chem, sorry! :/ but no, not at all. I can't grasp the concept one bit....

OpenStudy (whpalmer4):

if you decompose 1 mole of \(H_2O_2\), how many moles of oxygen result?

OpenStudy (whpalmer4):

and how much mass does one of those moles of oxygen have?

OpenStudy (tparker22):

2?

OpenStudy (tparker22):

I don't know :/

OpenStudy (whpalmer4):

okay, 2 is the right answer for the number of moles of oxygen.

OpenStudy (tparker22):

Ok

OpenStudy (whpalmer4):

the nifty thing about moles is that 1 mole of something has mass of whatever the atomic mass of 1 atom/molecule/whatever, in grams. So if you are talking about hydrogen, for example, which has an atomic mass of just a hair over 1, 1 mole of hydrogen atoms has a mass of just a hair over 1 gram

OpenStudy (whpalmer4):

so, look up the atomic mass of oxygen in your periodic table, what do you get?

OpenStudy (tparker22):

Okay, so how would I put that into an equation?

OpenStudy (tparker22):

15.99

OpenStudy (whpalmer4):

let's understand the concept first before worrying about the equation. yes, my table says 15.9994 grams for 1 mole of oxygen atoms. so, we know how many moles of oxygen (14.3), and we know the mass of 1 mole of oxygen, how might we find the mass of 14.3 moles of oxygen?

OpenStudy (whpalmer4):

\[\text{mass} = \frac{\text{grams}}{\text{mole}} * \text{moles}\]

OpenStudy (tparker22):

Uh, would we use division?

OpenStudy (whpalmer4):

I just gave you the schematic... you know grams/mole (atomic mass), and you know moles (number of moles of oxygen)...

OpenStudy (whpalmer4):

If I tell you that 1 pound of dog kibble costs $3.99, and I want you to buy 2 pounds, how would you find the cost?

OpenStudy (whpalmer4):

unit price * number of units, right? exactly the same deal here. \[\text{mass of oxygen} = \frac{15.9994\text{ grams oxygen}}{1\text{ mole oxygen}} * 14.3\text{ moles oxygen}\]\[\text{mass of oxygen} = \frac{15.9994\text{ grams oxygen}}{1\cancel{\text{ mole oxygen}}} * 14.3\cancel{\text{ moles oxygen}}\]\[\text{mass of oxygen} =15.9994\text{ grams oxygen} * 14.3 = \]

OpenStudy (whpalmer4):

but I've deliberately made a mistake here...do you see what it is?

OpenStudy (whpalmer4):

hint: didn't we decide that each mole of hydrogen peroxide decomposed into 2 moles of oxygen?